# RF-MatID: Dataset and Benchmark for Radio Frequency Material Identification

### v1

@@ 1 Introduction
Material identification is the task of identifying an object’s physical category (e.g., metal, plastic) from its intrinsic properties, which is fundamental for autonomous systems, particularly in embodied intelligence. Embodied intelligence arises from an agent’s physical interaction with the environment, enabling it to perceive, reason, and adapt. Accurate material identification allows an embodied agent to infer object properties within its operational range, guiding manipulation and functional interactions. This capability has broad implications across diverse domains, such as enabling embodied agents to interpret fine-grained object attributes for physical scene understanding through material perception (xiao2018unified), and to ground complex reasoning and control in material-driven functional affordances (do2018affordancenet)
@@
To overcome the inherent constraints of optical sensors, non-visual sensing modalities, such as radio frequency (RF), are gaining attention as they exploit electromagnetic interactions to reveal intrinsic material properties beyond surface appearance (khushaba2022radar; hagele2025occluded). However, RF-based material identification has not yet been extensively explored. There are no publicly accessible large-scale datasets, hindering reproducibility and fair benchmarking across algorithms. In addition, commercial off-the-shelf (COTS) sensors offer limited and fragmented frequency coverage, hindering systematic evaluation of material identification across the RF spectrum, which is crucial for selecting optimal operational bands in various applications. Finally, most studies are conducted in controlled laboratory settings and rarely incorporate perturbations such as variations in sensor–object geometry (e.g., incidence angle, stand-off-distance), leaving open questions regarding robustness and scalability in deployment scenarios.
Thus, we present RF-MatID, a large-scale, wide-band, and geometry-diverse RF dataset designed to advance fine-grained material identification and enable the development of more robust algorithms. RF-MatID encompasses 16 carefully curated fine-grained categories, systematically derived from 5 superclasses to capture subtle intra-class variations. The dataset spans a broad frequency spectrum from 4 to 43.5 GHz, sampled at 53 points per GHz. Through large-scale data acquisition, RF-MatID comprises 142k samples represented in both the frequency and time domains (71k samples in each). Note that these 142k samples correspond to representation-level instances derived from 71k unique physical measurements, each provided with a paired frequency- and time-domain representation. RF-MatID also provides 5 frequency protocols (including protocols compliant with legal frequency regulations in major global economies) and 7 split settings for a comprehensive benchmark. The key contributions and characteristics of RF-MatID are summarized as follows:
@@ 2.2 Radio-Frequency Material Datasets
Currently, most RF-based material datasets are collected using wireless COTS sensors operating in the 0.9–81 GHz frequency range, capturing electromagnetic signals that encode material properties (chen2024wireless). Accordingly, these datasets can be broadly categorized based on the sensing modality into four clusters. Radio-Frequency-Identification (RFID)-based datasets are mostly collected in the 0.90–0.927 GHz frequency band, and the target object is either tagged on the surface or placed between a reader and tags deployed in the environment (wang2017tagscan; ha2020food). The modality enables applications such as through-wall sensing via strong low-frequency penetration, but the reliance on physical tags greatly complicates data collection. Wi-Fi-based datasets primarily leverage the 2.4 GHz and 5 GHz frequency bands, and the target object is typically placed between a Wi-Fi transmitter and receiver (feng2019wimi; shi2021environment). The ubiquity of Wi-Fi devices facilitates convenient data collection, but the narrow bandwidth and distortions induced by hardware imperfections constrain the reliability of signals. Ultra-WideBand (UWB)-based datasets are mostly collected using ultra-wideband pulse signals in the 3.1–10.6 GHz frequency range and the equipment setup is similar to WiFi-based collection. UWB-based systems can capture fine-grained attenuation and delay properties (dhekne2018liquid; zheng2021siwa), but face practical barriers from strict synchronization demands and high deployment costs. Millimeter-Wave (MmWave)-based datasets are mainly collected using COTS devices operating in the 57–64 GHz and 76–81 GHz frequency bands (wu2020msense; shanbhag2023contactless; chen2025material). A single mmWave radar unit is an integrated transceiver that provides accurate phase information characterizing the dielectric properties of the materials. However, its limited penetration confines the sensing to material surfaces, making it highly sensitive to environmental occlusions.
@@ 4.2 Data collection and preprocessing
Concretely, for each material, samples are systematically collected across various distances and angles configurations. Data acquisition begins at a distance of 200 mm, with an angle of 0°, where 20 samples are recorded. Each sample consists of 2048 frequency-domain bins covering the 4–43.5 GHz band. The angle is then incremented from 1° to 10°, with 10 samples collected at each angle. After completing all angles at the current distance, the distance is increased in 50 mm steps up to 2000 mm, and the same angular sampling procedure is repeated at each step. The process ensures dense coverage of the distance–angle space while increasing the sampling number at normal incidence. Appendix Figure 6 & 7 shows the effect of distance and incidence angle on the frequency signal. Other forms of realistic perturbations are also discussed in the appendix 8.4.
@@ 4.3 Intended Uses
RF-MatID is designed to facilitate a broad spectrum of research and applications, especially in the field of embodied AI. As the first publicly available RF material identification dataset and benchmark with diverse protocols and settings, it enables fair comparisons and accelerates the advancement of material identification algorithms. By providing a standardized dataset aligned with calibrated radar systems, RF-MatID enables the evaluation of models’ transferability to real-world application scenarios. For instance, models can be trained on the 24 GHz sub-band in RF-MatID and then evaluated on radar signals captured by embodied agents operating in practical indoor scenarios, enabling assessment of their zero-shot or few-shot transferability. RF-MatID further facilitates research on domain adaptation and generalization through cross-domain evaluation settings, advancing algorithm robustness under diverse domains and environmental conditions. From a modality perspective, RF-MatID introduces a compact, low-cost sensing platform operating in the UWB–mmWave spectrum. The platform is specifically customized for indoor embodied AI applications, as it can be integrated onto a robot’s end-effector to enable fine-grained local material characterization. This capability supports material aware manipulation and affordance driven workflows, such as selecting grasp strategies according to material compliance, adjusting contact forces based on surface hardness, and enabling downstream policies that rely on material grounded affordance cues, for example graspable, cuttable, or pourable objects. Moreover, the sensing platform can serve as a complementary modality within multimodal learning frameworks, providing materials’ electromagnetic characteristics that enrich the information available for embodied perception. For example, when combined with vision-based perception pipelines in indoor scene understanding, the UWB-mmWave signals provide fine grained local material characteristics that can be further mapped to material grounded affordances of individual objects. When fused with visual cues that capture spatial structure and object geometry in cluttered environments, this enables embodied agents to achieve more comprehensive scene understanding across multiple dimensions, ultimately supporting more reliable reasoning and action in complex indoor settings.
@@ 5.1 Benchmark Setup
In our benchmark, Accuracy is reported in all experimental analyses. To address the limitation that accuracy may be dominated by majority classes, we also report the Macro F1-score alongside accuracy when presenting baseline results under different task configurations. We further include precision and recall, with the detailed results reported in the appendix 7.

### v2

@@ 1 Introduction
Material identification is the task of identifying an object’s physical category (e.g., metal, plastic) from its intrinsic properties, which is fundamental for physical artificial intelligence (Physical AI), particularly in embodied intelligence. Embodied intelligence arises from an agent’s physical interaction with the environment, enabling it to perceive, reason, and adapt. Accurate material identification allows an embodied agent to infer object properties within its operational range, guiding manipulation and functional interactions. This capability has broad implications across diverse domains, such as enabling embodied agents to interpret fine-grained object attributes for physical scene understanding through material perception (xiao2018unified), and to ground complex reasoning and control in material-driven functional affordances (do2018affordancenet)
@@
To overcome the inherent constraints of optical sensors, non-visual sensing modalities, such as radio frequency (RF), are gaining attention as they exploit electromagnetic interactions to reveal intrinsic material properties beyond surface appearance (khushaba2022radar; hagele2025occluded). However, RF-based material identification has not yet been extensively explored. There are no publicly accessible large-scale datasets, hindering reproducibility and fair benchmarking across algorithms. In addition, commercial off-the-shelf (COTS) sensors offer limited and fragmented frequency coverage (yang2024mm), hindering systematic evaluation of material identification across the RF spectrum, which is crucial for selecting optimal operational bands in various applications. Finally, most studies are conducted in controlled laboratory settings and rarely incorporate perturbations such as variations in sensor–object geometry (e.g., incidence angle, stand-off-distance), leaving open questions regarding robustness and scalability in deployment scenarios.
Thus, we present RF-MatID, a large-scale, wide-band, and geometry-diverse RF dataset designed to advance fine-grained material identification and enable the development of more robust Physical AI algorithms. RF-MatID encompasses 16 carefully curated fine-grained categories, systematically derived from 5 superclasses to capture subtle intra-class variations. The dataset spans a broad frequency spectrum from 4 to 43.5 GHz, sampled at 53 points per GHz. Through large-scale data acquisition, RF-MatID comprises 142k samples represented in both the frequency and time domains (71k samples in each). Note that these 142k samples correspond to representation-level instances derived from 71k unique physical measurements, each provided with a paired frequency- and time-domain representation. RF-MatID also provides 5 frequency protocols (including protocols compliant with legal frequency regulations in major global economies) and 7 split settings for a comprehensive benchmark. The key contributions and characteristics of RF-MatID are summarized as follows:
@@ 2.2 Radio-Frequency Material Datasets
Currently, most RF-based material datasets are collected using wireless COTS sensors operating in the 0.9–81 GHz frequency range, capturing electromagnetic signals that encode material properties (chen2024wireless). Accordingly, these datasets can be broadly categorized based on the sensing modality into four clusters. Radio-Frequency-Identification (RFID)-based datasets are mostly collected in the 0.90–0.927 GHz frequency band, and the target object is either tagged on the surface or placed between a reader and tags deployed in the environment (wang2017tagscan; ha2020food). The modality enables applications such as through-wall sensing via strong low-frequency penetration, but the reliance on physical tags greatly complicates data collection. Wi-Fi-based datasets primarily leverage the 2.4 GHz and 5 GHz frequency bands, and the target object is typically placed between a Wi-Fi transmitter and receiver (yang2023sensefi; feng2019wimi; shi2021environment). The ubiquity of Wi-Fi devices facilitates convenient data collection, but the narrow bandwidth and distortions induced by hardware imperfections constrain the reliability of signals. Ultra-WideBand (UWB)-based datasets are mostly collected using ultra-wideband pulse signals in the 3.1–10.6 GHz frequency range and the equipment setup is similar to WiFi-based collection. UWB-based systems can capture fine-grained attenuation and delay properties (dhekne2018liquid; zheng2021siwa), but face practical barriers from strict synchronization demands and high deployment costs. Millimeter-Wave (MmWave)-based datasets are mainly collected using COTS devices operating in the 57–64 GHz and 76–81 GHz frequency bands (wu2020msense; shanbhag2023contactless; chen2025material). A single mmWave radar unit is an integrated transceiver that provides accurate phase information characterizing the dielectric properties of the materials. However, its limited penetration confines the sensing to material surfaces, making it highly sensitive to environmental occlusions.
@@ 4.2 Data collection and preprocessing
Concretely, for each material, samples are systematically collected across various distances and angles configurations. Data acquisition begins at a distance of 200 mm, with an angle of 0°, where 20 samples are recorded. Each sample consists of 2048 frequency-domain bins covering the 4–43.5 GHz band. The angle is then incremented from 1° to 10°, with 10 samples collected at each angle. After completing all angles at the current distance, the distance is increased in 50 mm steps up to 2000 mm, and the same angular sampling procedure is repeated at each step. The process ensures dense coverage of the distance–angle space while increasing the sampling number at normal incidence. Appendix Figure 6 & 7 show the effect of distance and incidence angle on the frequency signal. Other forms of realistic perturbations are also discussed in the appendix 8.4.
@@ 4.3 Intended Uses
RF-MatID is designed to facilitate a broad spectrum of research and applications, especially in the field of embodied AI. As the first publicly available RF material identification dataset and benchmark with diverse protocols and settings, it enables fair comparisons and accelerates the advancement of material identification algorithms. By providing a standardized dataset aligned with calibrated radar systems, RF-MatID enables the evaluation of models’ transferability to real-world application scenarios. For instance, models can be trained on the 24 GHz sub-band in RF-MatID and then evaluated on radar signals captured by embodied agents operating in practical indoor scenarios, enabling assessment of their zero-shot or few-shot transferability. RF-MatID further facilitates research on domain adaptation and generalization through cross-domain evaluation settings, advancing algorithm robustness under diverse domains and environmental conditions. From a modality perspective, RF-MatID introduces a compact, low-cost sensing platform operating in the UWB–mmWave spectrum. The platform is specifically customized for indoor embodied AI applications, as it can be integrated onto a robot’s end-effector to enable fine-grained local material characterization. This capability supports material aware manipulation and affordance driven workflows, such as selecting grasp strategies according to material compliance, adjusting contact forces based on surface hardness, and enabling downstream policies that rely on material grounded affordance cues, for example graspable, cuttable, or pourable objects. Moreover, the sensing platform can serve as a complementary modality within multimodal learning frameworks, providing materials’ electromagnetic characteristics that enrich the information available for embodied perception (chen2025xfi; AN2026101429). For example, when combined with vision-based perception pipelines in indoor scene understanding, the UWB-mmWave signals provide fine grained local material characteristics that can be further mapped to material grounded affordances of individual objects. When fused with visual cues that capture spatial structure and object geometry in cluttered environments, this enables embodied agents to achieve comprehensive scene understanding across multiple dimensions, ultimately supporting more reliable reasoning and action in complex indoor settings.
@@ 5.1 Benchmark Setup
In our benchmark, Accuracy is reported in all experimental analyses. To address the limitation that accuracy may be dominated by majority classes, we also report the Macro F1-score alongside accuracy when presenting baseline results under different task configurations. We further include precision and recall, with the detailed results reported in the appendix Table 7.